Solve the system by substitution.
y=−4x
y=3x+28

Answer by mananth(16949)

Solve the system by substitution.
y=−4x
y=3x+28
substitute y in y=3x+28
-4x =3x+28
subtract 3x from both sides
-4x-3x = 3x-3x+28
-7x = 28
/-7

x = 28/-7
x=-4
y=-4x
y=(-4)(-4)=16
The solution
x = -4
y = 16